Find the equation of straight line which passes through point (3,-2) and cuts off positive interception on x axis and y axis which are in ratiob4:3

Grade:11

1 Answers

Arun
6 years ago
Dear student
Let the line:
xa + y/b = 1
Now
a/b = 4/3 
Hence
a = 4k, b = 3k
Now
3/4k - 2/3k = 1
(9-8)/ (12k) = 1
12 k = 1
k = 1/12
Hence a = 4*1/12 = 1/3
b = 3* 1/12 = ¼
Now equation of line:
3x + 4y = 1
Regards
Arun (askIITians forum expert)